# Lite-Mind: Towards Efficient and Robust Brain Representation Learning

### v2

@@ Abstract
Research in decoding visual information from the brain, particularly through the non-invasive fMRI method, is rapidly progressing. The challenge arises from the limited data availability and the low signal-to-noise ratio of fMRI signals, leading to a low-precision task of fMRI-to-image retrieval. State-of-the-art MindEye remarkably improves fMRI-to-image retrieval performance by leveraging a large model, i.e., a 996M MLP Backbone per subject, to align fMRI embeddings to the final hidden layer of CLIP’s Vision Transformer(ViT). However, significant individual variations exist among subjects, even within identical experimental setups, mandating the training of large subject-specific models. The substantial parameters pose significant challenges in deploying fMRI decoding on practical devices. To this end, we propose Lite-Mind, a lightweight, efficient, and robust brain representation learning paradigm based on Discrete Fourier Transform, that efficiently aligns fMRI voxels to fine-grained information of CLIP. Our experiments demonstrate that Lite-Mind achieves an impressive 94.6% fMRI-to-image retrieval accuracy on the NSD dataset for Subject 1, with 98.7% fewer parameters than MindEye. Lite-Mind is also proven to be able to be migrated to smaller fMRI datasets and establishes a new state-of-the-art for zero-shot classification on the GOD dataset. Code will be available at https://github.com/gongzix/Lite-Mind.
1 Introduction
@@
Brain decoding holds immense significance in elucidating intricate mental processes and serves as a pivotal technology for brain-computer interfaces [26, 14, 28]. Among diverse brain decoding tasks, the decoding of visual information stands out as a paramount yet challenging endeavor, allowing us to unravel the complex mechanisms involved in visual processing, object recognition, scene understanding, and even higher-order cognitive functions. In the pursuit of decoding natural visual information, functional magnetic resonance imaging (fMRI) is widely employed as a non-invasive modality to decipher the perceptual and semantic information intricately encoded within the cerebral cortex [22]. It draws considerable attention to fMRI-to-image retrieval/reconstruction tasks. This paper focuses primarily on tasks related to fMRI-to-image retrieval.
@@
Figure 1: The significance of lightweight brain decoding model. Subject-specific models are visually represented by clouds of distinct colors, and the cloud size correlates with model parameters.
@@
The success of fMRI-to-image retrieval heavily relies on aligning fMRI signals with the representation space of images. However, fMRI data often suffers from spatial redundancy, noise, and sample sparsity, leading to poor representation of fMRI signals and potential overfitting to noise distribution for deep models. To address these challenges, previous studies have found ridge regression and shallow linear models to be effective solutions for fMRI-based models [29, 15, 9, 1, 13, 48]. Recently, it has been demonstrated the effectiveness of leveraging pretrained CLIP models [30] as a powerful bridge between fMRI voxels and images, since CLIP’s image embeddings capture fine-grained and semantic information of pictures [23, 25, 21, 35]. Initially, the CLS embeddings of the CLIP image encoder was used to align voxels in Mind-Reader [21]. Considering the CLS embeddings alone provide limited information for fMRI data, BrainClip [23] and UniBrain [25] introduce the text annotations of images to assist in aligning fMRI voxels with image representations. Brain-diffuser [27] aligns voxel representations with the last hidden layer of CLIP using large-scale linear models ($257\times 14$M) for the first time.
@@
Among the CLIP-related models, MindEye [35] stands out by employing a large-scale (996M) MLP Backbone and Projector to map fMRI voxels and align their representations with the final hidden layer of CLIP using contrastive learning. MindEye achieves remarkable breakthroughs, surpassing $90\%$ accuracy for both image and brain retrieval, outperforming previous state-of-the-art accuracies that were below $50\%$. However, the practical application of large-scale models in brain science research faces limitations. Each subject’s fMRI data can exhibit significant variations, even within identical experimental setups, making it currently infeasible to establish a universal encoding model that performs well for all individuals. Consequently, training individual encoding models for each subject becomes necessary, as depicted in Figure 1. Apparently, it is impractical to expect hospitals to have access to computing resources with sufficient performance to train a unique large-scale model (e.g., MindEye) for each individual. Therefore, the ideal scenario involves equipping individuals with lightweight brain-computer interface edge devices, enabling image representations to be efficiently retrieved by a lightweight model. To this end, there is an urgent need to develop a brain network that is lightweight and efficient, as illustrated in the right segment of Figure 1, for practical implementation and deployment of brain-computer interfaces in real-world scenarios.
@@
MindEye strongly benefits from an MLP backbone with a high parameter count, which allows for straightforward and powerful compression and preservation. However, when dealing with fMRI data, which typically has a low signal-to-noise ratio, an MLP focusing on voxel value-wise mappings may inadvertently suppress informative voxel signal values while being inefficient at reducing noisy values. This can lead to model parameter redundancy and decelerated convergence speed. Recently, the Fourier Transform has gained attention in deep learning, demonstrating lightweight characteristics and improved efficiency in learning frequency domain patterns, due to its global perspective and energy concentration properties [47]. Additionally, the Fourier Transform naturally possesses the advantage of processing signals: it is more effective and efficient to filter out noise and maintain informative voxel signals in the frequency domain. It presents a promising alternative to large-scale MLPs for encoding fMRI signals.
@@
To tackle the trade-off between model capacity and scalability in fMRI-to-image retrieval, we propose Lite-Mind, a lightweight, efficient, and robust brain representation learning paradigm. We redesign MindEye with our elaborate Discrete Fourier Transform (DFT) Backbone, avoiding the large MLP Backbone used in MindEye. Extensive experiments show that Lite-Mind achieves 94.6% retrieval accuracy for Subject 1 on the NSD dataset, with 98.7% fewer parameters than MindEye. Lite-Mind also proves its adaptability to smaller brain datasets and establishes a new state-of-the-art zero-shot classification on the GOD dataset. Our contributions are summarized as follows:
@@
  • We demonstrated that the Fourier Transform has the advantage of lightweight and efficient for fMRI representation learning in frequency domain.
  • We proposed a new Brain Representation Network DFT Backbone based on Discrete Fourier Transform, as an efficient and robust fMRI representation network for fMRI-to-image retrieval. Its effectiveness, efficiency, and lightweight are demonstrated.
@@
2 Related Work
2.1 Brain Visual Decoding
@@
The study of visual decoding using fMRI in the human brain has been a long-standing endeavor. In the most important fMRI-image reconstruction and partial fMRI-image retrieval tasks, the foundation is how to align fMRI signals to the image representation space or intermediate space. Due to the low signal-to-noise ratio of fMRI, the initial research emphasized using models based on linear regression to extract fMRI signals and images into the intermediate space for functional decoding of the human brain or reconstruction of faces and natural scenes [38]. Further research utilizes pre-trained VGG to enrich hierarchical image features [13, 36, 7]. With the development of cross-modal tasks, the image-text pre-training model CLIP has been introduced into fMRI-image research. Mind-reader [21] is the first to adopt a contrastive learning approach, using the nsdgenal ROI region of NSD data and a shallow resnet-like model to align fMRI with the CLS embeddings output by the CLIP image encoder. Mind-Reader did not achieve good performance in fMRI-image retrieval when using InfoNCE loss. BrainClip [23] uses a VAE model for retrieval tasks while still using the CLS embeddings. Brain diffuser [27] introduced a fine-grained representation of the image of the last hidden layer in CLIP. But they used 257 different linear regression models to align each layer of the 257 hidden layers, resulting in the model being too cumbersome. Mind-Vis [5] conducted self-supervised mask pre-training on fMRI signals for another dataset, BOLD5000, and demonstrated that voxel patches can be used to process fMRI voxel values. MindEye [35] uses a large MLP backbone for contrastive learning, aligning the last hidden layer $257\times 768$ of CLIP, and using diffusion prior in DALLE·2 [31] for the first time to narrow the disjointed vectors. At the same time, it creatively proposed a mixed contrast loss MixCo as a data augmentation method to make the large model converge. The recent trend in brain fMRI decoding research tends to use larger models to achieve better performance on downstream tasks [42, 16], ignoring the privacy and deployment issues discussed above. However, to the best of our knowledge, no previous research has focused on lightweight and efficient brain networks.
2.2 Fourier Transform in Deep Learning
Fourier Transform plays a vital role in the area of digital signal processing. It has been introduced to deep learning for enhanced learning performance [11, 8, 45], and has demonstrated excellent performance in many fields, such as computer vision, natural language processing, and time series analysis. In the field of computer vision, GFNet [32] utilizes Fast Fourier Transform(FFT) to convert images to the frequency domain and exchange global information between learnable filters. As a continuous global convolution independent of input resolution, Guibas et al. [12] design the Adaptive Fourier Neural Operator(AFNO) frame token mixing. Xu et al. [43] devise a learning-based frequency selection method to identify trivial frequency components and improve image classification accuracy. As for natural language processing, Lee-Thorp et al. [20] use the Fourier Transform as a text token mixing mechanism on text classification. In research of time series forecast, Fourier Transform has deeper applications  [3, 18, 17, 44, 46, 19]. To increase the accuracy of multivariate time-series forecasting, Cao et al. [3] propose a spectral temporal graph neural network (StemGNN), which mines the correlations and time dependencies between sequences in the frequency domain. Yang et al. [44] propose bilinear temporal spectral fusion (BTSF), which updates the feature representation in a fused manner by explicitly encoding time-frequency pairs and using two aggregation modules: spectrum-to-time and time-to-spectrum. Yi et al. [47] prove that frequency-domain MLP is more efficient than time-domain MLP. Processing fMRI frequency domain information is more helpful for understanding the mechanism of collaborative operation within the human brain. However, there is currently no research on processing fMRI signals in the frequency domain for representation learning.
@@ 2.2 Fourier Transform in Deep Learning
Figure 2: Overview of our Lite-Mind. Figure (a) is the architecture of the MLP Backbone of MindEye. Figure (b) is the architecture of the DFT backbone and Retrieval Pipeline. fMRI voxels are inputted into DFT Backbone to obtain voxel embeddings for downstream tasks.
3 Lite-Mind
3.1 Overview
@@
As illustrated in Figure 2, Lite-Mind comprises two main components: 1) DFT Backbone consists of fMRI Spectrum Compression and fMRI Token Decoupling in Frequency-domain, mapping flattened voxels to an intermediate space. 2) Retrieval pipeline contains Diffusion Projector, Contrastive Learning, and Downstream tasks with the voxel embeddings from the DFT backbone.
3.2 DFT Backbone
@@
3.2.1 fMRI Spectrum Compression
Patchify and Embedding The fMRI data $x$ is one-dimensional data composed of voxels. we firstly divide $x$ into $n$ non-overlapping patches $x=[x_{1},x_{2},...,x_{n}]$ using zero-padding, and adopt convolutions with positional encoding to obtain multiple tokens $t=[t_{1},t_{2},...,t_{n}]$.
@@
After obtaining the spectrum of voxel tokens, we introduce learnable filters for filtering processing. Voxel spatial features are effectively consolidated within each frequency element, enabling the extraction of informative features from voxels through the point-wise product in the frequency domain. We use $\textbf{K}=[\textbf{k}_{1},\textbf{k}_{2},...,\textbf{k}_{M}]$ to represent the filter library, where $M$ is the number of filters in the filter library. Spectrum $\hat{X}$ after filtering and compression is obtained as follows:
$\hat{X}=\sum_{m=1}^{M}\frac{1}{n}{|X|}^{2}\odot\textbf{k}_{m}cos(\frac{(2m-1)\pi}{2M})$ (2)
where $\odot$ is the element-wise multiplication, $|X|^{2}$ is the power spectrum of $X$. The $|X|^{2}$ operation smooths the spectrum, highlighting the main components of the fMRI spectrum from a global perspective. It also facilitates voxel information concentration and noise filtering. $cos((2m-1)\pi/2M)$ adopts the filterling form of Discrete Cosine Transform (DCT), which compacts better energy and can aggregate the more important information in fMRI voxels. Its combination of application with the channel filter library K allows for efficient frequency domain feature compression and noise denoising for fMRI.
Finally, we employ Inverse Discrete Fourier Transform (IDFT), denoted by $F^{-1}(\cdot)$, to convert the spectrum back into the spatial domain:
$\hat{t}\leftarrow F^{-1}(\hat{X})$ (3)
3.2.2 fMRI Token Decoupling in Frequency-domain
To align voxel tokens to image tokens, we decouple the filtered and compressed fMRI tokens again in frequency domain. Similarly, we follow Equation (2) to convert $\hat{t}$ to the frequency domain, then decouple tokens by a MLP-like model in frequency domain(FreMLP).
FreMLP. Following Equation (2), the complex number output of $\hat{t}$ after DFT is $\hat{X}\in\mathbb{C}^{n\times d}$. Given a complex number weight matrix $\mathcal{W}\in\mathbb{C}^{n\times n^{\prime}}$, and a complex number bias $\mathcal{B}\in\mathbb{C}^{n^{\prime}}$, then the frequency-domain MLPs for token decoupling can be formulated as:
$X^{\prime}={\sigma(\hat{X}^{T}\mathcal{W}+\mathcal{B})}^{T}$ (4)

### v3

@@ Abstract.
The limited data availability and the low signal-to-noise ratio of fMRI signals lead to the challenging task of fMRI-to-image retrieval. State-of-the-art MindEye remarkably improves fMRI-to-image retrieval performance by leveraging a large model, i.e., a 996M MLP Backbone per subject, to align fMRI embeddings to the final hidden layer of CLIP’s Vision Transformer (ViT). However, significant individual variations exist among subjects, even under identical experimental setups, mandating the training of large subject-specific models. The substantial parameters pose significant challenges in deploying fMRI decoding on practical devices. To this end, we propose Lite-Mind, a lightweight, efficient, and robust brain representation learning paradigm based on Discrete Fourier Transform (DFT), which efficiently aligns fMRI voxels to fine-grained information of CLIP. We elaborately design a DFT backbone with Spectrum Compression and Frequency Projector modules to learn informative and robust voxel embeddings. Our experiments demonstrate that Lite-Mind achieves an impressive 94.6% fMRI-to-image retrieval accuracy on the NSD dataset for Subject 1, with 98.7% fewer parameters than MindEye. Lite-Mind is also proven to be able to be migrated to smaller fMRI datasets and establishes a new state-of-the-art for zero-shot classification on the GOD dataset.
1. Introduction
@@
Brain decoding holds immense significance in elucidating intricate mental processes and serves as a pivotal technology for brain-computer interfaces (Naselaris et al., 2011; Kamitani and Tong, 2005; Palazzo et al., 2020). Among diverse brain decoding tasks, the decoding of visual information stands out as a paramount yet challenging endeavor, allowing us to unravel the complex mechanisms involved in visual processing, object recognition, scene understanding, and even higher-order cognitive functions. In the pursuit of decoding natural visual information, functional magnetic resonance imaging (fMRI) is widely employed as a non-invasive modality to decipher the perceptual and semantic information intricately encoded within the cerebral cortex (Linden, 2021). It draws considerable attention to fMRI-to-image retrieval/reconstruction tasks. This paper focuses primarily on tasks related to fMRI-to-image retrieval.
@@
Figure 1. The significance of lightweight brain decoding model. Subject-specific models are visually represented by clouds of distinct colors, and the cloud size correlates with model parameters.
@@
The success of fMRI-to-image retrieval heavily relies on aligning fMRI signals with the representation space of images. However, fMRI data often suffers from spatial redundancy, noise, and sample sparsity, leading to poor representation of fMRI signals and potential overfitting to noise distribution for deep models. To address these challenges, previous studies have found ridge regression and shallow linear models to be effective solutions for fMRI-based models (Pereira et al., 2018; Kewenig et al., 2023; Emmanuele et al., 2021; Allen et al., 2022; Horikawa and Kamitani, 2017; Zou et al., 2022). Recently, it has been demonstrated the effectiveness of leveraging pretrained CLIP models (Radford et al., 2021) as a powerful bridge between fMRI voxels and images, since CLIP’s image embeddings capture fine-grained and semantic information of pictures (Liu et al., 2023; Mai and Zhang, 2023; Lin et al., 2022; Scotti et al., 2023). Initially, the CLS embeddings of the CLIP image encoder was used to align voxels in Mind-Reader (Lin et al., 2022). Considering the CLS embeddings alone provide limited information for fMRI data, BrainClip (Liu et al., 2023) and UniBrain (Mai and Zhang, 2023) introduce the text annotations of images to assist in aligning fMRI voxels with image representations. Brain-diffuser (Ozcelik and VanRullen, 2023) aligns voxel representations with the last hidden layer of CLIP using large-scale linear models ($257\times 14$M) for the first time.
@@
Among the CLIP-related models, MindEye (Scotti et al., 2023) stands out by employing a large-scale (996M) MLP Backbone and Projector to map fMRI voxels and align their representations with the final hidden layer of CLIP using contrastive learning. MindEye achieves remarkable breakthroughs, surpassing $90\%$ accuracy for both image and brain retrieval, outperforming previous state-of-the-art accuracies that were below $50\%$. However, the practical application of large-scale models in brain science research faces limitations. Each subject’s fMRI data can exhibit significant variations, even within identical experimental setups, making it currently infeasible to establish a universal encoding model that performs well for all individuals. Consequently, training individual encoding models for each subject becomes necessary, as depicted in Figure 1. Apparently, it is impractical to expect hospitals to have access to computing resources with sufficient performance to train a unique large-scale model (e.g., MindEye) for each individual. Therefore, the ideal scenario involves equipping individuals with lightweight brain-computer interface edge devices, enabling image representations to be efficiently retrieved by a lightweight model. To this end, there is an urgent need to develop a brain network that is lightweight and efficient, as illustrated in the right segment of Figure 1, for practical implementation and deployment of brain-computer interfaces in real-world scenarios.
@@
MindEye strongly benefits from an MLP backbone with a high parameter count, which allows for straightforward and powerful compression and preservation. However, when dealing with fMRI data, which typically has a low signal-to-noise ratio, an MLP focusing on voxel value-wise mappings may inadvertently suppress informative voxel signal values while being inefficient at reducing noisy values. This can lead to model parameter redundancy and decelerated convergence speed. Recently, the Fourier Transform has gained attention in deep learning, demonstrating lightweight characteristics and improved efficiency in learning frequency domain patterns, due to its global perspective and energy concentration properties (Yi et al., 2024). Additionally, the Fourier Transform naturally possesses the advantage of processing signals: it is more effective and efficient to filter out noise and maintain informative voxel signals in the frequency domain. It presents a promising alternative to large-scale MLPs for encoding fMRI signals.
@@
In order to mitigate the noise effect of fMRI and obtain more robust latent representations, we designed global filter-based Filter Blocks for fMRI denoising and fMRI spectrum compression. In addition to the global view property of frequency domain filtering, which can block out the noise and thus obtain better robust representations, the frequency domain computation also brings a great efficiency improvement compared to the large MLP structure. Since the features of the compressed fMRI we obtain are more concentrated in the frequency domain, we also design a Frequency Projector based on MLP in the frequency domain. Combining all our computational paradigms in frequency domain, we propose Lite-Mind, a lightweight, efficient, and robust brain representation learning paradigm. We redesign MindEye with our elaborate Discrete Fourier Transform (DFT) Backbone, avoiding the large MLP Backbone used in MindEye. Extensive experiments show that Lite-Mind achieves 94.6% retrieval accuracy for Subject 1 on the NSD dataset, with 98.7% fewer parameters than MindEye. Lite-Mind also proves its adaptability to smaller brain datasets and establishes a new state-of-the-art zero-shot classification on the GOD dataset. Our contributions are summarized as follows:
@@
  • We demonstrated that the Fourier Transform has the advantage of being lightweight and efficient for fMRI representation learning in the frequency domain.
  • We proposed a novel DFT Backbone with elaborate Spectrum Compression and Frequency Projector modules, which is theoretically and empirically verified lightweight, efficient, and robust for brain representation.
@@
2. Related Work
2.1. Brain Visual Decoding
@@
The study of visual decoding using fMRI in the human brain has been a long-standing endeavor. In the most important fMRI-image reconstruction and partial fMRI-image retrieval tasks, the foundation is how to align fMRI signals to the image representation space or intermediate space. Due to the low signal-to-noise ratio of fMRI, the initial research emphasized using models based on linear regression to extract fMRI signals and images into the intermediate space for functional decoding of the human brain or reconstruction of faces and natural scenes (Takagi and Nishimoto, 2023). Further research utilizes pre-trained VGG to enrich hierarchical image features (Horikawa and Kamitani, 2017; Shen et al., 2019; Du et al., 2023). With the development of cross-modal tasks, the image-text pre-training model CLIP has been introduced into fMRI-image research. Mind-reader (Lin et al., 2022) is the first to adopt a contrastive learning approach, using the nsdgenal ROI region of NSD data and a shallow Resnet-like model to align fMRI with the CLS embeddings output by the CLIP image encoder. Mind-Reader did not achieve good performance in fMRI-image retrieval when using InfoNCE loss. BrainClip (Liu et al., 2023) uses a VAE model for retrieval tasks while still using the CLS embeddings. Brain diffuser (Ozcelik and VanRullen, 2023) introduced a fine-grained representation of the image of the last hidden layer in CLIP. However they used 257 different linear regression models to align each layer of the 257 hidden layers, resulting in the model being too cumbersome. Mind-Vis (Chen et al., 2023) conducted self-supervised mask pre-training on fMRI signals for another dataset, BOLD5000, and demonstrated that voxel patches can be used to process fMRI voxel values. MindEye (Scotti et al., 2023) uses a large MLP backbone for contrastive learning, aligning the last hidden layer $257\times 768$ of CLIP, and using diffusion prior in DALLE·2 (Ramesh et al., 2022) for the first time to narrow the disjointed vectors. At the same time, it creatively proposed a mixed contrast loss MixCo as a data augmentation method to make the large model converge. The recent trend in brain fMRI decoding research tends to use larger models to achieve better performance on downstream tasks (Xia et al., 2024; Kneeland et al., 2023), ignoring the privacy and deployment issues discussed above. However, to the best of our knowledge, no previous research has focused on lightweight and efficient brain networks.
@@
Figure 2. Overview of our Lite-Mind. Figures (a) and (b) show the architecture of the MLP Backbone of MindEye and the DFT backbone and Retrieval Pipeline, respectively. fMRI voxels are inputted into DFT Backbone to obtain voxel embeddings.
2.2. Fourier Transform in Deep Learning
Fourier Transform plays a vital role in the area of digital signal processing. It has been introduced to deep learning for enhanced learning performance (Gueguen et al., 2018; Ehrlich and Davis, 2019; Yang and Soatto, 2020), and has demonstrated excellent performance in many fields, such as computer vision, natural language processing, and time series analysis. In the field of computer vision, GFNet (Rao et al., 2021) utilizes Fast Fourier Transform(FFT) to convert images to the frequency domain and exchange global information between learnable filters. As a continuous global convolution independent of input resolution, Guibas et al. (Guibas et al., 2021) design the Adaptive Fourier Neural Operator(AFNO) frame token mixing. Xu et al. (Xu et al., 2020) devise a learning-based frequency selection method to identify trivial frequency components and improve image classification accuracy. As for natural language processing, Lee-Thorp et al. (Lee-Thorp et al., 2021) use the Fourier Transform as a text token mixing mechanism for text classification. In research of time series forecast, Fourier Transform has deeper applications  (Cao et al., 2020; Lange et al., 2021; Koç and Koç, 2022; Yang and Hong, 2022). To increase the accuracy of multivariate time-series forecasting, Cao et al. (Cao et al., 2020) propose a spectral temporal graph neural network (StemGNN), which mines the correlations and time dependencies between sequences in the frequency domain. Yang et al. (Yang and Hong, 2022) propose bilinear temporal spectral fusion (BTSF), which updates the feature representation in a fused manner by explicitly encoding time-frequency pairs and using two aggregation modules: spectrum-to-time and time-to-spectrum. Yi et al. (Yi et al., 2024) prove that frequency-domain MLP is more efficient than time-domain MLP. Processing fMRI frequency domain information is more helpful for understanding the mechanism of collaborative operation within the human brain. However, there is currently no research on processing fMRI signals in the frequency domain for representation learning.
3. Lite-Mind
3.1. Overview
@@
As illustrated in Figure 2, Lite-Mind comprises two main components: 1) DFT Backbone consists of fMRI Spectrum Compression and Frequency Projector, mapping flattened voxels to an intermediate space. 2) Retrieval pipeline contains Diffusion Projector, Contrastive Learning, and Downstream tasks with the voxel embeddings from the DFT backbone.
3.2. DFT Backbone
@@
3.2.1. fMRI Spectrum Compression
Patchify and Embedding The fMRI data $x$ after ROI preprocessing and flattening is a one-dimensional long vector composed of voxels. We conduct patchify firstly because it has been verified as simple and effective for spatial/temporal representation of long sequences or high-dimensional vectors. At first, we divide $x$ into $n$ non-overlapping patches $x=[x_{1},x_{2},...,x_{n}]$ using zero-padding, and adopt convolutions with positional encoding to obtain multiple tokens $t=[t_{1},t_{2},...,t_{n}]$.
@@
Voxel spatial features are effectively consolidated within each element of the frequency spectrum of voxel tokens, enabling the extraction of informative features from voxels through the point-wise product in the frequency domain. Accordingly, we introduce learnable filters $\textbf{K}=[\textbf{k}_{1},\textbf{k}_{2},...,\textbf{k}_{M}]$ to filter and compress the spectrum as follows, where $\mathbf{K}$ denotes the filter library and $M$ is the number of filters in the library.
(2) $\hat{X}=\sum_{m=1}^{M}\frac{1}{n}{|X|}^{2}\odot\textbf{k}_{m}cos(\frac{(2m-1)\pi}{2M})$
where $\odot$ is the element-wise multiplication, $|X|^{2}$ is the power spectrum of $X$. The $|X|^{2}$ operation smooths the spectrum, highlighting the main components of the fMRI spectrum from a global perspective. $cos((2m-1)\pi/2M)$ adopts the filtering form of Discrete Cosine Transform (DCT), which compacts better energy and can aggregate the significant information in fMRI voxels. Its combination with the channel filter library K allows for efficient feature compression and noise denoising for fMRI from the frequency domain.
Finally, we employ Inverse Discrete Fourier Transform (IDFT) $F^{-1}(\cdot)$ to convert the spectrum back into the spatial domain:
(3) $\hat{t}\leftarrow F^{-1}(\hat{X})$
3.2.2. Frequency Projector
The fMRI spectrum after filtering has a more concentrated feature in the frequency domain compared to the spatial domain, which is caused by the energy concentration effect of the frequency domain calculation. We therefore designed a Frequency Projector in the frequency domain to align the image tokens, which has the added benefit of maintaining all computations of our DFT Backbone in the frequency domain, where lightness and efficiency are guaranteed. To align voxel tokens to image tokens, we mapped the filtered and compressed fMRI tokens by an MLP-like projector in the frequency domain(FreMLP). Similarly, we follow Equation (2) to convert $\hat{t}$ to the frequency domain.
FreMLP. Following Equation (2), the complex number output of $\hat{t}$ after DFT is $\hat{X}\in\mathbb{C}^{n\times d}$. Given a complex number weight matrix $\mathcal{W}\in\mathbb{C}^{n\times n^{\prime}}$, and a complex number bias $\mathcal{B}\in\mathbb{C}^{n^{\prime}}$, then the FreMLP can be formulated as:
(4) $X^{\prime}={\sigma(\hat{X}^{T}\mathcal{W}+\mathcal{B})}^{T}$